# Large-scale Long-tailed Disease Diagnosis on Radiology Images

### v2

@@ Abstract
In this study, we aim to investigate the problem of large-scale, large-vocabulary disease classification for radiologic images, which can be formulated as a multi-modal, multi-anatomy, multi-label, long-tailed classification. Our main contributions are three folds: (i), on dataset construction, we build up an academically accessible, large-scale diagnostic dataset that encompasses 5568 disorders linked with 930 unique ICD-10-CM codes, containing 39,026 cases (192,675 scans). (ii), on model design, we present a novel architecture that enables to process arbitrary number of input scans, from various imaging modalities, which is trained with knowledge enhancement to leverage the rich domain knowledge; (iii), on evaluation, we initialize a new benchmark for multi-modal multi-anatomy long-tailed diagnosis. Our method shows superior results on it. Additionally, our final model serves as a pre-trained model, and can be finetuned to benefit diagnosis on various external datasets.
@@ 1 Introduction
In the ever-evolving landscape of clinical medicine, the advent of radiology techniques, such as X-ray, CT, MRI, and ultrasound, has truly revolutionized the medical field, offering a non-invasive yet deeply revealing perspective of the human body for disease diagnosis and management. These imaging techniques are now at the cusp of a new era with the integration of artificial intelligence (AI).
Recent literature highlights the significant potential of developing diagnostic models in the medical field, the developments can be generally cast into two categories: one is a specialist, which has already shown success in identifying and managing a wide array of diseases [15, 78, 86]. However, a notable limitation of these models is their specialization, as they often focus on a narrow range of disease categories, targeting limited anatomical regions, based on specific imaging modalities. This specialization restricts their ability to fully address the diverse and complex cases encountered in real-world clinical settings.
While on the other extreme, there is an emerging trend towards developing Generalist Medical Artificial Intelligence (GMAI) models [56, 81, 85]. Taking inspiration from the breakthroughs in natural language processing and computer vision, these models aim to amalgamate data from diverse sources, including various imaging methods, patient histories, and current medical research, to offer more comprehensive diagnostic and treatment solutions. Nonetheless, the development of GMAI models faces formidable challenges, such as the need for substantial computational power, meticulously curated multimodal datasets covering an extensive range of medical conditions and patient demographics, and advanced models equipped to tackle unique medical intricacies, like extremely unbalanced data distribution and the need for domain-specific expertise.
In this paper, we consider the problem of large-scale, large-vocabulary disease classification for radiologic images, marking a transition phase between specialist and generalist models. Specifically, compared to existing specialists, we aim to initiate the research for developing a computational model that can handle multi-modal, multi-anatomy, and multi-label disease diagnosis, in the face of extremely unbalanced distribution, embracing a wider scope of diagnoses across various anatomical regions and imaging modalities. In contrast to generalist models, our investigation offers a more feasible and targeted playground for exploring sophisticated algorithms in academic labs, offering opportunities for detailed error analysis, which is often impractical in the development of large-scale generalist models due to prohibitive computational costs. Overall, we make contributions from three aspects, namely, a large-scale open dataset and its construction pipeline, preliminary model architecture exploration, and an evaluation benchmark.
On dataset construction, we build up an academically accessible, large-scale diagnostic dataset derived from Radiopaedia [5]. Each sample is associated to the International Classification of Diseases, i.e., ICD-10-CM [1], indicating the diagnostic category, for example, ‘S86’ refers to ‘Injury of muscle, fascia and tendon at lower leg level’. The dataset naturally displays an unbalanced distribution, varying from 1 to 964 cases in each disease category. Additionally, each case within the dataset involves multiple multi-modal scans, with the number of modalities ranging from 1 to 5, the number of images per case spanning between 1 and 30. As a result, we construct a long-tailed, multi-scan medical disease classification dataset, as shown in Figure 1, with 39,026 cases (192,675 scans) across 7 human anatomy regions and 9 diverse modalities covering 930 ICD-10-CM codes, 5568 disorders [^1] Disorders encompass a range of conditions such as abnormality, syndromes, injuries, poisonings, signs, symptoms, findings and diseases. Diseases are specific pathological conditions characterized by a set of identifiable signs and symptoms., termed as Radiopaedia3D Diagnosis Dataset (RP3D-DiagDS). We will release all the data, complete with corresponding disorders, ICD-10-CM codes, and detailed definitions.
On architecture design, we demonstrate a new model that supports both 2D and 3D input from various modalities, together with a transformer-based fusion module for comprehensive diagnosis. Specifically, for visual encoding, we explore two variants of backbones, namely, ResNet-based, and ViT-based to perform 2D or 3D unified encoding. Then we fuse the multi-scan information with a transformer-based fusion module, treating each image embedding as an input token. At training time, we adopt knowledge-enhanced training strategy [79, 84], specifically, we leverage the rich domain knowledge to pre-train a knowledge encoder with natural language and use it to guide the visual representation learning for disease diagnosis.
On evaluation, we carry out a series of ablation studies on the effectiveness of different training configurations, for example, visual backbones (ViT or ResNet), augmentation implementation, and depth of 3D input volumes. Then, we evaluate the model on our proposed benchmark of multi-modal multi-scan long-tailed multi-label diagnosis and demonstrate the superiority of our proposed methods on it. Furthermore, our trained model showed strong transferring abilities. By fine-tuning, it can benefit numerous external datasets, regardless of their image dimensions, imaging modalities, and shooting anatomies.
@@
2 Related Work
2.1 Disease Classification Datasets
Open-source datasets play a crucial role in the development of AI for medical image analysis. Unlike natural images, the curation of medical image datasets presents unique challenges due to factors, such as privacy concerns, requirement of specialized domain knowledge for annotation, etc. In the literature, there has been a number of open-source datasets for disease diagnosis. Notably, large-scale datasets such as NIH ChestX-ray [76], MIMIC-CXR [38], and CheXpert [34] stand out for their comprehensive collection of annotated X-ray images, facilitating extensive research and advancements in automated disease classification. However, it’s important to note three key limitations. First, these open-source diagnosis datasets mainly consist of consist of chest X-rays [63, 12, 76, 34, 50, 66, 31, 36, 71, 25, 59, 22, 38, 19], thus are 2D images. There are only a few 3D datasets available, with a limited number of volumes [54, 3, 2, 4, 39, 43, 51, 33, 58, 66, 10, 7, 9]. Second, a significant portion of these large-scale datasets focuses solely on binary classification of specific diseases [71, 25, 31, 36]. Third, existing large-scale datasets containing a broad range of disease categories in different granularities [12, 82, 53], for example, in PadChest [12], disorders exist at different levels, including infiltrates, interstitial patterns, and reticular interstitial patterns. Such variation in granularity poses a great challenge for representation learning. In summary, these datasets fall short of meeting real-world clinical needs, which often involve complex, multimodal, and multi-image data from single patient. Therefore, constructing a dataset that mirrors the intricacies of actual clinical scenarios is necessary.
2.2 Specialist Diagnosis Models
The prevailing paradigm in earlier diagnostic models is a specialized model trained on a limited range of disease categories. These models focus on specific imaging modalities and are targeted towards particular anatomical regions. Specifically, ConvNets are widely used in medical image classification due to its outstanding performance, for example, [21, 69, 73, 29, 19, 46, 13, 74] have demonstrated excellent results on identifying a wide range of diseases. Recently, Vision Transformer (ViT) has garnered immense interest in the medical imaging community, numerous innovative approaches  [20, 80, 27, 70, 52, 87, 83, 65, 84, 67, 79, 48, 72, 26, 55] have emerged, leveraging ViTs as a foundation for further advancements in this field.
2.3 Generalist Medical Foundation Models
The other stream of work is generalist medical foundation models  [75, 81, 56, 85]. These models represent a paradigm shift in medical AI, aiming to create versatile, comprehensive AI systems capable of handling a wide range of tasks across different medical modalities, by leveraging large-scale, diverse medical data. MedPaLM M [75] reaches performance competitive with or exceeding the state-of-the-art (SOTA) on various medical benchmarks, demonstrating the potential of the generalist foundation model in disease diagnosis and beyond. RadFM [81] is the first medical foundation model capable of processing 3D multi-image inputs, demonstrating the versatility and adaptability of generalist models in processing complex imaging data. However, the development of GMAI models requires substantial computational power, which is often impractical for exploring sophisticated algorithms in academic labs.
2.4 Long-tailed Classification Strategy
Medical image diagnosis inherently faces long-tail challenges, as the prevalence of common diseases is usually substantially higher than that of rare diseases. A straightforward solution to the imbalance problem is to perform re-sampling during training [24, 28, 42, 49], e.g., over-sampling on tailed classes or under-sampling on head classes. However, this approach often triggers over-fitting in tail classes, resulting insufficient training for the head classes. Loss re-weighting is another widely used solution to tackle the long-tailed distribution problem [47, 68]. Focal loss [45, 47] refines cross-entropy by assigning lower weights to easily learnable data, thereby prioritizing challenging or misclassifiable instances. These approaches have primarily been explored on relatively small datasets with a limited number of categories. In this paper, our objective is to initiate the problem for large-scale, long-tailed, multi-scan medical disease classification.
3 Dataset Construction
In this section, we present the details of our dataset, RP3D-DiagDS, that follows a similar procedure as RP3D [81]. Specifically, cases in our dataset are sourced from the Radiopaedia website [5] – a growing peer-reviewed educational radiology resource website, that allows the clinicians to upload 3D volumes to better reflect real clinical scenarios. Additionally, all privacy issues have already been resolved by the clinicians at uploading time. It is worth noting that, unlike RP3D [81] that contains paired free-form text description and radiology scans for visual-language representation learning, here, we focus on multi-modal, multi-anatomy, and multi-label disease diagnosis (classification) under extremely unbalanced distribution.
Overall, the proposed dataset contains 39,026 cases, of 192,675 images from 9 diverse imaging modalities and 7 human anatomy regions, note that, each case may contain images of multiple scans. The data covers 5,568 different disorders, that have been manually mapped into 930 ICD-10-CM [1] codes.
3.1 Data Collection
Here, we describe the procedure of our data curation process, shown in Figure 2. Specifically, we collect three main components from each case on Radiopaedia webpage, namely, ‘Patient data’, ‘Radiology images’, and ‘Articles’  (example webpage shown in Figure 2 (a)). ‘Patient data’ includes brief information of this patient, for example, age, gender, etc. ‘Radiology images’ denote a series of radiology examination scans. Figure 5 (a) provides a statistical analysis on the number of images within one single case. ’Articles’ contains links to related articles named with corresponding disorders, which are treated as diagnosis labels and have been meticulously peer-reviewed by experts in Radiopaedia Editorial Board [^2] https://radiopaedia.org/editors.
Figure 2: The data collection pipeline. In (a) we show two original cases. Our dataset presents a multi-scan multi-modal multi-label long-tailed diagnosis task with 2D and 3D scan involed. In (b) we show the general data curation progress. We totally collect 39026 cases with 192675 images labeled by 5568 disorder classes and 930 ICD-10-CM classes. In addition to the intermediate case number, we also show the intermediate image number together with the final normal and abnormal case number in (b).
To start with, we can collect 50,970 cases linked with 10,670 articles from Radiopaedia website. However, naively adopting the article titles as disorders may lead to ambiguities from two aspects: (i) not all articles are related to disorders, (ii) article titles can be written with different granularity, for example, “pneumonia” and “bacterial pneumonia” can be referred to as different disorders, though they should ideally be arranged in a hierarchical structure, (iii) normal cases from Radiopaedia are not balanced in modalities and anatomies. Next, we discuss 3-stage procedure to alleviate the above-mentioned challenges.
Article Filtering. We leverage the GPT-4 [64] to automatically filter the article list, leaving those refer to disorders. Specifically, taking inspiration from the self-consistency prompts [77], we design two different query prompts with similar meanings, as shown in Figure 3 (a). An article name is labeled as disorder if GPT-4 gives consistent positive results from both prompts, while for those GPT-4 gives inconsistent results, we do manually check them. To measure the quality for filtering, we randomly sample a portion of data for manual checking to control its quality. The confusion metrics are shown in Figure 3 (c). The $100\%$ precision score indicates our auto-filtering strategy, can strictly ensure the left ones to be disorders. Eventually, 5342 articles can pass the first auto-criterion and 226 pass the second manually checking, resulting in 5,568 disorder classes, as shown in Figure 2 (b). Cases not linked to any articles are excluded, ultimately yielding 38,858 cases.
Figure 3: The article filtering pipeline. In (a), we show the two GPT prompts and (b) is the whole pipeline where we leverage the two prompts to query GPT-4 twice and only the consistent results will be used for auto-filtering, otherwise, we will check it manually. (c) shows the accuracy of the auto-filtering cases. Note that, all the 200 article samples evaluated in (c) have a consistent prediction under the two GPT-4 prompts.
Mapping disorders to ICD-10-CM. Upon getting disorder classes, they may fall into varying hierarchical granularity levels, here, we hope to map them to internationally recognized standards. For this purpose, we utilize the International Classification of Diseases, Tenth Revision, Clinical Modification (ICD-10-CM) codes, which is a customized version of the ICD-10 used for coding diagnoses in the U.S. healthcare system. The ICD was originally designed as a health care classification system, providing a system of diagnostic codes for classifying diseases, including nuanced classifications of a wide variety of signs, symptoms, abnormal findings, complaints, social circumstances, and external causes of injury or disease. Specifically, we hire ten medical PhD students to manually map the article titles into ICD-10-CM codes. After cross-checking by a ten-year clinician, we have mapped the 5,568 disorders into the corresponding ICD-10-CM codes. We unify various disorders into the first hierarchical level of ICD-10-CM code tree, resulting in 930 ICD-10-CM classes. For example, both “J12.0 Adenoviral pneumonia” and “J12 Viral pneumonia” are mapped to the code “J12 Viral pneumonia, not elsewhere classified”, dismissing the ambiguities caused by diagnosis granularity. Note that, we will provide the ICD-10-CM codes along with the original 5,568 classes, each accompanied by its corresponding definition, thus, the dataset can be employed for training diagnosis or visual-language models.
Adding Normal Cases. Despite there are normal cases on Radiopaedia [^3] https://radiopaedia.org/articles/normal-imaging-examples, covering most of anatomies and modalities, we additionally collect more normal cases from MISTR [^4] https://mistr.usask.ca/odin/, with images available for research under a CC-BY-NC-SA 4.0 license. The expanded normal cases include 168 cases. The distribution of modalities and anatomical regions will be shown in the data statistics section later.
Finally, we get 39,026 cases containing 192,675 images labeled by 5,568 disorder classes and 930 ICD-10-CM classes and will continually maintain the dataset, growing the case number.
3.2 Dataset Statistics
In this section, we provide detailed analysis on our proposed dataset, from three aspects, namely, modality coverage, anatomy coverage, and disease coverage.
Table 1: Comparison between RP3D-DiagDS and existing disease diagnosis datasets. #Image refers to the number of images in the dataset. #Disorders represents the number of categories of labeled disorders. 2D/3D indicates the dimension of images in the dataset, and the case-level column indicates whether the dataset contains multiple images per patient. Our dataset demonstrates significant increase in both size and diversity.
Dataset Anatomy Modality #Image #Disorders 2D 3D Case-level
Brain-Tumor [3] Head and Neck MRI 7k 4
CE-MRI  [17] Head and Neck MRI 3k 3
Brain-Tumor-17  [4] Head and Neck MRI 4.4k 17
OASIS[39, 43, 51] Head and Neck MRI 3.2k 2
ICH2020 [33] Head and Neck CT 82 5
COVID19-CT-DB [58] Chest CT 1k 9
Covid-CXR2 [66] Chest X-ray 19k 3
POCUS [10] Chest Ultrasound 1.1k 3
MIA-COV19 [40] Chest CT 5.3k 2
LIDC-IDRI [7] Chest CT 1.0k 3
NIH Chestxray [76] Chest X-ray 100k 14
PadChest [12] Chest X-ray 160k 193
CheXpert [34] Chest X-ray 224k 14
RSNA [71] Chest X-ray 30k 2
SIIM-ACR [25] Chest X-ray 15k 2
MIMIC-CXR [38] Chest X-ray 371k 14
VinDr-CXR [59] Chest X-ray 18k 28
VinDr-PCXR [63] Chest X-ray 9k 51
VinDr-Mammo [61] Breast Mammography 20k 10
DDSM [32, 44] Breast Mammography 55k 5
CMMD2022 [14] Breast Mammography 5.2k 2
BUSI [6] Breast Ultrasound 780 3
VinDr-SpineXr [62] Spine X-ray 10k 13
MRNet [9] Knee MRI 1370 3
DeepLesion [82] Whole body Radiology 32k 22
RadImageNet [53] Whole body CT, MRI, Ultrasound 1.35M 165
RP3D-DiagDS (ours) Whole body Radiology 192k 5568 (ICD-10-CM:930)*
    We map disorders to ICD-10-CM for standardization.
Analysis on Modality Coverage. RP3D-DiagDS comprises images from 9 modalities, namely, computed tomography (CT), magnetic resonance imaging (MRI), X-ray, Ultrasound, Fluoroscopy, Nuclear medicine, Mammography, DSA (angiography), and Barium Enema. Each case may include images from multiple modalities, to ensure precise and comprehensive diagnosis of disorders. Overall, approximately 19.4% of the cases comprise images from two modalities, while around 2.9% involve images from three to five modalities. The remaining cases are associated with image scans from a single modality. The distribution of modalities among all abnormal samples is illustrated in Figure 4 (a). The modalities of normal cases follow similar distributions with the abnormal cases, as shown in Figure 4 (b).
(a) Abnormal Cases Distribution in Modality
(b) Normal Cases Distribution in Modality
(c) Abnormal Cases Distribution in Anatomy
(d) Normal Cases Distribution in Anatomy
Figure 4: The distribution of imaging modalities of abnormal (a) and normal (b) cases in RP3D-DiagDS and anatomies of abnormal (c) and normal (d) cases in RP3D-DiagDS. Each label is annotated with the class name, number of cases, and the corresponding proportion.
Analysis on Anatomy Coverage. RP3D-DiagDS comprises images from various anatomical regions, including the head and neck, spine, chest, breast, abdomen and pelvis, upper limb, and lower limb, providing comprehensive coverage of the entire human body. The statistics are shown in Figure 4 (c) and 4 (d).
Figure 5: Distribution Analysis. In (a), we show the distribution for the number of images in one case. Each case may include multiple images from patient history scans, different modalities, different angles or conditions. In (b), we demonstrate the long-tailed distributions for disorder and ICD-10-CM classes. We also categorize these classes into three categories: “head class”, “body class” and “tail class” based on the number of cases. Notably, to better show the main part of the case distributions in (b), we clip the axes, indicated by the dotted axes lines.
Analysis on Disease Coverage. For both disorder and disease classification, each case can correspond to multiple disorders, resulting in RP3D-DiagDS a multi-label classification dataset. As shown in Figure 5 (b), the distributions exhibit extremely long-tailed pattern, rendering such 2D & 3D image classification problem as a long-tailed multi-label classification task. We define the ‘head class’ category with case counts greater than 100, the ‘body class’ category with case counts between 30 and 100, and the ‘tail class’ category with case counts less than 30.
4 Method
In this section, we aim to initiate a preliminary investigation on computational architectures for large-scale, long-tailed disease diagnosis on radiology, specifically, we start by defining the problem of case-level multi-label classification in Sec. 4.1, then we elaborate the architecture in Sec. 4.2, and knowledge-enhanced training strategy in Sec. 4.3.
4.1 Problem Formulation
We consider one case with its disease labels, represented as $\mathcal{X}=\{x_{1},\dots,x_{S},y_{1},\dots,y_{c}\}$, where $x_{i}$ denotes a specific scan (2D or 3D) under a certain radiologic examination, for example, MRI, CT or X-ray within $\mathcal{X}$, $S$ denotes the scan number for the case and $c$ represents the total diagnostic labels. Note that, there may be multiple scans from the same or different modalities for each patient.

### v3

@@ Abstract
Developing a generalist radiology diagnosis system can greatly enhance clinical diagnostics. In this paper, we introduce RadDiag, a foundational model supporting 2D and 3D inputs across various modalities and anatomies, using a transformer-based fusion module for comprehensive disease diagnosis. Due to patient privacy concerns and the lack of large-scale radiology diagnosis datasets, we utilize high-quality, clinician-reviewed radiological images available online with diagnosis labels. Our dataset, RP3D-DiagDS, contains 40,936 cases with 195,010 scans covering 5,568 disorders (930 unique ICD-10-CM codes). Experimentally, our RadDiag achieves 95.14% AUC on internal evaluation with the knowledge-enhancement strategy. Additionally, RadDiag can be zero-shot applied or fine-tuned to external diagnosis datasets sourced from various hospitals, demonstrating state-of-the-art results. In conclusion, we show that publicly shared medical data on the Internet is a tremendous and valuable resource that can potentially support building a generalist AI for healthcare.
@@ 1 INTRODUCTION
In the landscape of clinical medicine, the emergence of radiology techniques, for example, X-ray, CT, and MRI, has truly revolutionized the medical field, which enables to diagnose by reading radiological images. Recent literature has demonstrated significant potential for developing AI models for disease diagnosis in the medical field [83, 72]. However, a notable limitation of the existing models is their specialization, as these ‘specialist’ models often focus on a narrow range of disease categories, and thus usually fail to accommodate the breadth and complexity of clinical scenarios. Conversely, an ideal radiology diagnosis system is expected to comprehensively analyze the combination of arbitrary historical scans across any anatomical regions and imaging modalities for a wide range of diseases.
In this paper, we aim to build a foundational radiology diagnostic model, that supports information fusion from an arbitrary number of 2D and 3D inputs across different modalities and anatomic sites. Generally speaking, there are three major challenges. First, lack of datasets for training. Despite the large volume of radiology data generated per day across the world, publicly available ones are extremely limited due to privacy issues. Existing open-source datasets are normally collected from a certain modality or anatomy with very limited diseases, thus significantly hindering the development of powerful disease diagnosis models. Second, lack of unified architecture formulation. Models for diagnosing radiological diseases are often tailored to particular datasets, which are optimized for a narrow spectrum of diseases linked with particular imaging modalities [22, 34]. In contrast, developing a foundation model for disease diagnosis requires the architecture to be capable of fusing information from an arbitrary number of images of various modalities, for example, X-ray, CT, and MRI. Third, lack of benchmark to monitor progress. Benchmarking the models’ performance on disease diagnosis relies on datasets that well simulate real-world scenarios, for instance, containing many diseases, long-tail distribution, etc. Existing benchmarks typically focus on a very limited number of diseases, with single modality inputs [11, 51]. To comprehensively measure the progress in disease diagnosis, a high-quality benchmark has yet to be established.
@@
To overcome the above-mentioned challenges, we consider to exploit the widely available data on Internet, especially those focusing on high-quality radiology samples from professional institutions and educational websites, that have undergone meticulous de-identification, and reviewed by a panel of experienced clinicians. Specifically, we develop a flexible data collection pipeline, with each sample associated with an International Classification of Diseases code, i.e., ICD-10-CM [1], indicating the diagnostic category. The dataset presents a long-tailed distribution across disease categories. The case numbers range from 1 to 964 with some rare diseases. It includes multi-modal scans per case, with modalities varying from 1 to 5 and image counts between 1 and 30. As a result, we collect a multi-scan medical disease classification dataset, with 40,936 cases (195,010 scans) across 7 anatomy regions and 9 modalities, covering 930 ICD-10-CM codes and 5568 disorders, including a broad spectrum of medical conditions.
Architecturally, we conduct preliminary explorations on the design that supports both 2D and 3D input from various modalities, together with a transformer-based fusion module for comprehensive diagnosis. Specifically, for visual encoding, we explore three variants of backbones, namely, ResNet-based, ViT-based and ResNet-ViT-mixing to perform 2D or 3D unified encoding. Then, we fuse the multi-scan information with a transformer-based fusion module, treating each image embedding as an input token. At training time, we adopt knowledge-enhanced training strategy [77, 83], where we first leverage the rich domain knowledge to pre-train a knowledge encoder with natural language, and use it to guide the visual representation learning for disease diagnosis.
In experiments, we evaluate our model on a benchmark, constructed by images from online resources, featuring large-scale, multi-modal, multi-scan, long-tailed disease diagnosis. Additionally, we demonstrate that our proposed model can function as a general-purpose image encoder to capture a better image representation for radiology, making it possible to train and classify new sets via data-efficient fine-tuning or zero-shot transferring, leading to improved performance across diverse disease diagnosis tasks on numerous datasets, regardless of their image dimensions, modalities and anatomies. This paper demonstrates that publicly shared medical information is a tremendous resource that can be harnessed to advance the development of general AI for healthcare.
2 RESULTS
In this section, we start by briefly introducing our considered problem formulation, followed with showing the characteristics of our dataset for case-level multi-modal, multi-scan, long-tailed disorder/disease diagnosis. Then, we leverage a test split of the dataset to carry out a series of studies on diagnostic performance and establish a case-level diagnosis benchmark. Lastly, we train and assess the developed model accordingly, along with its transferring ability to various external datasets.
Problem Formulation
In this paper, our objective is to develop a radiology diagnostic model capable of integrating information from various images within a single case. We define a case as encompassing a collection of both 2D and 3D radiological images from different modalities and anatomical sites of a patient, who may be diagnosed with multiple conditions.
More specifically, we represent a case and its associated disease labels as $\mathcal{X}=\{x_{1},\dots,x_{S},y_{1},\dots,y_{c}\}$, where each $x_{i}$ signifies a specific scan (either 2D or 3D) from a radiological examination such as MRI, CT, or X-ray. Here, $S$ is the number of scans for the case, and $c$ is the total number of diagnostic labels. It is noteworthy that a patient’s case might include multiple scans across the same or different modalities.